A pump is lifting 200L of water per minute through a height of 5 m and ejects it through a nozzle of 2 cm in diameter. What is the velocity of water ejected and power of the pump?

Grade:

1 Answers

Vikas TU
5 years ago
Dear Sudent,
From the continuity equation Area x velocity = constant.
Therefore,
Rate  =  AV
1 m^3 = 1000 L
Then in 200 L = 0.2 m^3
0.2m^3/60s = 3.14*((0.02)^2/4)*v
 
0.0033333 = 0.00314*v
v = 1.0615 m/s